APCO Company manufectures vanious tnes of bicycles. Because of the high volurne of each tne, the company employs a process costing system using the weighted-average method. Bicycle parts are manufactured in the molding department and then are consolidated Into a single bike unit in the molding department and transferred to the assembly department where they are assembled After assembly. the bicycle is sent to the packing department. Annual cost and production agures for the assembly department are presented in the schedules below are identified at the inspection point when the ossembly labor processs 70% complete all assembly meteres hve been added at this point The norrnal rejection rate for detectve booes is 5% of the beyoes reading me inspection port Any defective bicycles above and destroyed the 5% quota are considered to be abnormal All defective bikes are removed Som me proooon process Assembly Department Coat Data Transferred-in Total Cost of from Molding Assembly Assembly Bike through Department Materials Conversion Ansembly S 106,070 1, 579, 270 s252, 120一了1,685,340 Prior pexiod eostn Current period $86,200 S 7,260 12,610 1,241,000 98,760239,510 106,020 Assenbly Department Production Data Percentage Complete Assenbly Asambly Bicycles Transferred-in Materials-Comme 3,040 Tzansferred-in from molding during year 49,000 Transterzed out to packing during year , 000 1001 100 100 100 100% 100 50 200 20 Ending inventory 4,080 Compute the number of defective, or spoiled, bikes that are considered to be a. Normal spoilage. b. Abnormal spoilage. Normal spoilege Abnormal spoilege
Answer #1
Normal spoilage 2246
Abnormal spoilage 3714

Working:

Physical Units
Beginning inventory 3040
Transferred-in from molding 49000
Total units to account for 52040
Transferred out to packing 42000
Ending inventory 4080
Spoilage 5960
Total units accounted for 52040

Normal spoilage = 5% x (49000 - 4080) = 5% x 44920 = 2246

Abnormal spoilage = 5960 - 2246 = 3714

Note: Normal spoilage is not computed on the beginning inventory which is 80% complete with respect to Assembly conversion since the inspection on same would already have been completed and spoilage identified when it reached 70% completion with respect to Assembly conversion. Of the 49000 units transferred-in from molding, 4080 units are in ending inventory which are less than 70% complete and hence not subject to inspection yet. Hence, the normal spoilage is computed on 49000 - 4080 = 44920 units.